OpenStudy (anonymous):

i need help for this equation y + 2 = -2(x + 3).

OpenStudy (anonymous):

pless help

OpenStudy (luigi0210):

What do you need help with? Putting it in slope form?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

First distribute the parenthesis

OpenStudy (anonymous):

y+2=2x+6

OpenStudy (luigi0210):

Wait, did you forget it's -2?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

\[-1*1=-1\] \[-1*-1=1\]

OpenStudy (anonymous):

y+2=-2x-6

OpenStudy (luigi0210):

There you go, now move that 2.

OpenStudy (anonymous):

where i dont get that part

OpenStudy (luigi0210):

Subtract 2 from both sides.

OpenStudy (anonymous):

y+2x+4

OpenStudy (luigi0210):

Close, what's -6-2

OpenStudy (anonymous):

8

OpenStudy (luigi0210):

Yup, don't forget the negative though: So it's y=-2x-8

OpenStudy (anonymous):

do u divide

OpenStudy (luigi0210):

Nope, you're done, that's your final answer

OpenStudy (luigi0210):

For slope-intercept form at least.

OpenStudy (anonymous):

subtract 2 from both sides first. and you have solved for y:)

OpenStudy (anonymous):

how about coordinates

OpenStudy (luigi0210):

Coordinates?

OpenStudy (anonymous):

on a graph

OpenStudy (luigi0210):

Oh you want to graph it too?

OpenStudy (anonymous):

yes

OpenStudy (luigi0210):

I recommend finding the x and y intercepts first

OpenStudy (anonymous):

never mind got it

OpenStudy (luigi0210):

Alright, good luck.